OpenStudy (anonymous):

Solve the quadratic equation by completing the square. –3x^2 + 9x = 1

OpenStudy (anonymous):

Okay, -3x^2 + 9x -1 =0

OpenStudy (anonymous):

Okay, I'm sorry. I just re-read about completing the square and the first step is to move the constant to the other side of the equation. So the original equation is where we start.

OpenStudy (anonymous):

From the original equation first we will divide by -3

OpenStudy (anonymous):

okay, ..

OpenStudy (anonymous):

x^2 - 3x = -1/3

OpenStudy (anonymous):

okay, now what do you do

OpenStudy (anonymous):

we take half of the coefficient of the x term and square it, adding it to both sides of the equation. So... (-3/2)^2 = 9/4 x^2 - 3x + 9/4 = -1/3 + 9/4

OpenStudy (anonymous):

I am literally just following these steps http://www.purplemath.com/modules/sqrquad.htm

OpenStudy (anonymous):

now we have a factorable left side. (x-3/2)^2 = 23/12

OpenStudy (anonymous):

x-3/2 = +/- sqrt 23/12

OpenStudy (anonymous):

x= 3/2 +/- sqrt 23/12

OpenStudy (anonymous):

{(3^1/2)x - 3(3^1/2)}^2=23/4

OpenStudy (anonymous):

typing eq on ths is a tough job